OpenStudy (anonymous):

How do you find the derivative of (x+4)^2?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Use chain rule. So lets say u=x+4 then the derivative of u^2 => 2u and the derivative is of u is 1 so you get 2*(x+4) *1

OpenStudy (anonymous):

frank gonna tell u

OpenStudy (anonymous):

It comes from \[dy/dx = dx/du * du/dy\]

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Chain_rule This is the wiki page concerning the chain rule

OpenStudy (mayankdevnani):

use (a+b)^2=a^2+b^2+2ab

OpenStudy (anonymous):

Thank you Mayan I forgot about that way to do it.